what is the longest side of a triangle with a perimeter of 39 if a similar triangle has side lengths of 3,4, and 6
1)12
2)13
3)18
4)20

Answer by Alan3354(69443)

3 + 4 + 6 = 13
39/13 = 3
3*6 = 18, the longest side